The current gain for a transistor working as common-base amplifier is 0.96. If the emitter current is 7.2 mA, then the base current is


        (a)    0.29mA


        (b)    0.35mA


        (c)    0.39mA


(d)    0.43mA


Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(a)

Current gain (a) =0.96

        Ic = 0.96 × 7.2 mA = 6.91 mA

        Ic = Ic + Ib

     Ib = Ie – Ic = 7.2 - 6.91 = 0.29 mA